# Large Language Models and Impossible Language Acquisition: “False Promise” or an Overturn of our Current Perspective towards AI

### v1

@@ 4.2.1 Setup
Our first experiment was conducted on a smaller dataset, which includes 10,000 simple SVO sentences, generated by a python program that is provided our GitHub repository, thus eliminate the potential inference of context, allowing us to establish a corpus adapting to Chomsky’s CFG. Appendix A includes details on preprocessing and formatting.
@@ 4.2.3 Results
There are clear distinctions between those are trained in natural and in impossible language, where Welch’s t-test indicates that the loss for parity negation group is particularly higher than control group ($p<.001$, t(305.0)=19.66t(305.0)=-19.66, Cohen’s d=2.20d=-2.20). The overturn of our primary hypothesis indicates that GPT-2 small model is better at learning languages that corresponds to the Chomskyan definition of natural languages, which leads to an inversion of the first hypothesis in the following experiment.
@@ 4.3.3 Results
Results from experiment 2 reveals a more nuanced pattern. The difference in loss values between natural and parity negation group, although apparently exist, was not statistically significant ($p=.223$, $t(799.9)=1.22$), whereas such difference do exist between natural and reversed group ($p<.001$, Cohen’s d=2.59d=-2.59), suggesting that loss may not thoroughly capture GPT2’s structural bias under larger dataset.
The hypothesis is validated. Integrating with the gaps of both loss values and perplexities in the second round, GPT-2 small model is inferred to have a preference in possible languages.
@@ 4.6 Further Conjecture
A phenomenon appears in the experiments could be relevant to the discoveries in McCoy et al. (2020), a paper finding that the inductive bias of a neural network is not neutral, but heavily impacted by the architecture of the network. Specifically, only models with the tree-structures have developed the bias for hierarchical syntactic rules, whereas sequential models failed to do so even with hints.
We notice that the GPT2 small models trained on impossible languages actually have better performance in negation parity group (a linear transformation) than in the reverses group (a more thorough transformation, destroying all the serial order), showing identity with the aforementioned findings. Such correlation could be seen as a kind of gradient sensitivity to the degree of preservation of syntactic linearity, which is necessarily a phenomenon under the influence of attention mechanism in transformer architecture.
@@ 5 Discussion and Conclusions
Our findings indicate an intrinsic preference for natural languages among GPT-2 small models, empirically falsify Chomsky’s criticism of ChatGPT. We have found numerous pieces of evidence to back up the argument, including lower loss values and perplexities in the control group compared to that in experimental groups, where models are forced to learn impossible languages.
@@
In summary, our work suggests that the “false promise” of ChatGPT is, in fact, an origin for a new understanding of intelligence. We are look forward to a future of AI beyond the constraints of metaphysical debates on whether machines possess “true understanding,” but a future where developing methodologies that empirically describe and assess the functionality of intelligent behavior. By synthesizing Piagetian constructivism and Rylean logical behaviorism in contemporary machine learning, we have pointed to a new interdisciplinary paradigm.
@@ Appendix A Preprocessing
In our first experiment, after generating the original sentences, we used another python script to transform them into two impossible language datasets with the spacy library. Sentences are firstly splitted, and then dealt with.
In our second experiment, as mentioned before, we employed a BabyLM dataset that includes about 100 million words with the aim to approximate the extent of lexicon that a child before the age of 12 could be exposed to (Gilkerson et al., 2017). Another python script is specified to execute the impossible language transformation. We still used the space library, but had to split the complete text into smaller chunks at the beginning. The whole process took several hours in total.

### v4

@@ 4.2.1 Setup
Our first experiment was conducted on a smaller dataset, which includes 10,000 simple SVO sentences, generated by a python program that is provided in our GitHub repository, thus eliminating the potential inference of context, allowing us to establish a corpus adapting to Chomsky’s CFG. Appendix A includes details on preprocessing and formatting.
@@ 4.2.3 Results
There are clear distinctions between those that are trained in natural and in impossible language, where Welch’s t-test indicates that the loss for parity negation group is particularly higher than control group ($p<.001$, t(305.0)=19.66t(305.0)=-19.66, Cohen’s d=2.20d=-2.20). The overturn of our primary hypothesis indicates that GPT-2 small model is better at learning languages that corresponds to the Chomskyan definition of natural languages, which leads to an inversion of the first hypothesis in the following experiment.
@@ 4.3.3 Results
Results from experiment 2 reveal a more nuanced pattern. The difference in loss values between the natural and parity negation groups, although apparently existing, was not statistically significant ($p=.223$, $t(799.9)=1.22$), whereas such difference does exist between the natural and reversed groups ($p<.001$, Cohen’s d=2.59d=-2.59), suggesting that loss may not thoroughly capture GPT2’s structural bias under larger dataset.
The hypothesis is validated. Integrating with the gaps of both loss values and perplexities in the second round, GPT-2 small model is inferred to have a preference for possible languages.
@@ 4.6 Further Conjecture
A phenomenon appearing in the experiments could be relevant to the discoveries in McCoy et al. (2020), a paper finding that the inductive bias of a neural network is not neutral, but heavily impacted by the architecture of the network. Specifically, only models with the tree-structures have developed the bias for hierarchical syntactic rules, whereas sequential models failed to do so even with hints.
We notice that the GPT2 small models trained on impossible languages actually have better performance in negation parity group (a linear transformation) than in the reverses group (a more thorough transformation, destroying all the serial order), showing consistency with the aforementioned findings. Such correlation could be seen as a kind of gradient sensitivity to the degree of preservation of syntactic linearity, which is necessarily a phenomenon under the influence of attention mechanism in transformer architecture.
@@ 5 Discussion and Conclusions
Our findings indicate an intrinsic preference for natural languages among GPT-2 small models, empirically falsifying Chomsky’s criticism of ChatGPT. We have found numerous pieces of evidence to back up the argument, including lower loss values and perplexities in the control group compared to that in experimental groups, where models are forced to learn impossible languages.
@@
In summary, our work suggests that the “false promise” of ChatGPT is, in fact, an origin for a new understanding of intelligence. We look forward to a future of AI beyond the constraints of metaphysical debates on whether machines possess “true understanding,” toward a future where developing methodologies that empirically describe and assess the functionality of intelligent behavior. By synthesizing Piagetian constructivism and Rylean logical behaviorism in contemporary machine learning, we have pointed to a new interdisciplinary paradigm.
@@ Appendix A Preprocessing
In our first experiment, after generating the original sentences, we used another python script to transform them into two impossible language datasets with the spaCy library. Sentences are first split, and then dealt with.
In our second experiment, as mentioned before, we employed a BabyLM dataset that includes about 100 million words with the aim to approximate the extent of lexicon that a child before the age of 12 could be exposed to (Gilkerson et al., 2017). Another python script is specified to execute the impossible language transformation. We still used the spaCy library, but had to split the complete text into smaller chunks at the beginning. The whole process took several hours in total.